OpenStudy (anonymous):

find the 6th term of the sequence 8,32,128

OpenStudy (paxpolaris):

its a geometric sequence with common ratio: 4

OpenStudy (anonymous):

is that the answer? or do i put it in the geometric sequence formula?

Parth (parthkohli):

\( \color{Black}{\Rightarrow g_n = g_1 \times r^{n - 1}}\)

OpenStudy (paxpolaris):

put it in that geometric sequence formula

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8 \times 4^5}\)

OpenStudy (anonymous):

I got 32768 is that right?

Parth (parthkohli):

4^5 = 1024 It'd be 1024 * 8

OpenStudy (anonymous):

oh i just got my math wrong thanks

Parth (parthkohli):

:)